OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1= 4 and a6= 24? 1) 128 2) 132 3) 136 4) 140

OpenStudy (anonymous):

i won't answer

OpenStudy (anonymous):

marc this is wrong

OpenStudy (dumbcow):

angela don't answer

OpenStudy (angela210793):

Marc...here are the formulas...please USE them Arithmetic prog: a2-a1=d an=an-1+d an=a1+(n-1)*d Sn=(a1+an)*n/2

OpenStudy (anonymous):

why?

OpenStudy (angela210793):

I wasn't...i got tired...I'm just giving the formulas

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I have no clue how to use that formula though

OpenStudy (angela210793):

Geometric prog:a2/a1=r an=an-1*r an=a1*r^(n-1) Sn=a1*(r^n -1)/(r-1)

OpenStudy (anonymous):

128

OpenStudy (angela210793):

all u have to do is replace...(and take a look at ur math book)

OpenStudy (anonymous):

Can you show me step by step with this problem what you mean by replace so i can see it better

OpenStudy (angela210793):

u have a1 and a6 and u have to find a32,a32 from the formula would be a32=a1+(n-1)*d u don't have d so u have to find it from wht ur given a6 from the formula is a6=a1+(n-1)*d 24=4+5d 5d=20 d=4(I'm supposing u know how to solve this) a32 from the formula would be a32=a1+(n-1)*d=4+(32-1)*4=4+31*4=128

OpenStudy (anonymous):

a1=a Therefore a2=a+d and a3=a+2d a4=a+3d... therefore a6=a+5d

OpenStudy (anonymous):

Find a and d, Then find a32=a+31d

OpenStudy (anonymous):

a1=4 therefore a = 4 a6=a+5d therefore a+5d=24 but a=4 so 4+5d=24 therefore d=4 so to find a32, a32=a+31d where a=4 and d=4 so a32= 4 + 31X4 which equals whatever

OpenStudy (angela210793):

which equals 128 :)